OpenStudy (anonymous):

f''(x) = 6x + sinx , Find f

OpenStudy (anonymous):

i dont know what to do with the 6x

OpenStudy (zzr0ck3r):

what is the derivative of 3x^2?

OpenStudy (anonymous):

do we not work with antiderivatives?

OpenStudy (zzr0ck3r):

you do, the derivative os 3x^2 = 6x so the anti derivative of 6x = 3x^2

OpenStudy (anonymous):

3x^2 is an antiderivative of 6x

OpenStudy (zzr0ck3r):

the anti derivative of a*x^n = a/(n+1) * x^(n+1)

OpenStudy (zzr0ck3r):

so do the opposite, increment the exponent by 1, then devide the coefficent by that number

OpenStudy (anonymous):

so the answer is x3-sinx+c ?

OpenStudy (anonymous):

x^3 - sinx +c

OpenStudy (anonymous):

Is the problem given the first or second derivative? I can't see....

OpenStudy (anonymous):

it's the second

OpenStudy (anonymous):

Oh.. ok.. then i shouldn't have erased... :(

OpenStudy (anonymous):

The first time you took the anti derivative you'll get the first derivative f'. So you should have gotten f'(x) = 3x^2 - cosx + C1. To get f, you'll have to take the anti derivative again... So f(x) = x^3/6 - sinx + C1x + C2 Try taking the derivative twice to see that you have the correct second derivativ.... \(\large f(x)=\frac{x^3}{6}-sinx+C_1x+C_2 \) where C1 and C2 are constants.

OpenStudy (anonymous):

isn't the antiderivative of 3x^2 = 3x^3/3 = x^3?

OpenStudy (anonymous):

Yes it is...

OpenStudy (anonymous):

so where did the x^3/6 derive from?

OpenStudy (anonymous):

Oh... lemme check...

OpenStudy (anonymous):

\(\large f(x)=x^3-sinx+C_1x+C_2 \) you're right... :) I think I'm too stuffed with turkey....

OpenStudy (anonymous):

haha Happy Thanksgiving!

OpenStudy (anonymous):

Same to u and everyone else!